# The Prospect from the Upcoming CMB Experiment LiteBIRD to Discover Axion-like Particles Using Milky Way

### v1

@@ 1 Introduction
Axions and axion-like particles (ALPs) are hypothetical particles predicted by the Beyond Standard Model (BSM) theories [23, 24, 25, 26, 27]. Axions were introduced as a solution to the strong-CP problem [28]. Axions are produced as a result of the global Peccei-Quinn U(1) symmetry breaking by the vacuum expectation value of a scalar field. Since axions and ALPs possess mass, they are Pseudo Nambu Goldstone Bosons (PNGBs) and make one of the dark matter candidates alongside Primordial Black Holes (PBHs), MAssive Compact Halo Objects (MACHOs), Weakly Interacting Massive Particles (WIMPs), sterile neutrinos, etc., and may only weakly interact with photons [23, 24, 25, 26, 27, 29, 28, 30, 31].
@@
Such a conversion can also take place in the Milky Way, as the photons travel through the ionized medium of the galactic halo in the presence of magnetic fields of the order of $\mu$G [51, 52, 53, 13]. This secondary anisotropy also leads to spectral distortions in the CMB. The spatial and spectral behaviour of the ALP signal is very different from that of CMB, and can be used to probe the signal [45, 46, 54, 47, 55]. Thus, based on the magnetic fields and electron densities observed in the Milky Way, accompanied by the spectral behaviour of the ALP signal, ALPs of mass range $10^{-15}$ to $10^{-11}$ eV can be probed. For this analysis, we consider the phenomenon of CMB photon-ALP resonant conversion in the Milky Way [43]. and show how the measurements can be improved over the current bounds from particle physics experiment CAST (CERN Axion Solar Telescope) at $g_{a\gamma}<6.6\times 10^{-11}\,\mathrm{GeV}^{-1}$ [56].
@@ 2.2 The spatial distribution of the electron density and magnetic field in the Milky Way
The $\mathrm{sech^{2}}$ form is used as it facilitates a rounder variation of the electron density at $z=0$, as compared to an exponential dependence. The $U(r)$ denotes the step function ($=1$ for $r>0$, and $=0$ for $r<0$), and the values of the parameters for the two components are shown in Table 0(b).
@@ 2.3 Simulating the ALP signal in the Milky Way
Using the formalism describe above, we can model the photon to ALPs conversion signal for different ALP masses and coupling strengths. We show a few representative simulations of the polarization maps for ALP masses $5\times 10^{-14}$ eV, $10^{-14}$ eV and $5\times 10^{-15}$ eV using the mentioned magnetic field and electron density models and coordinate transforming them to spherical coordinates, with the solar system as the origin. The observation point is the solar system, which is at the distance of about 8.5 kpc from the galactic center. To obtain the signal from different directions at this observation point, we coordinate transform to the solar system frame which is the origin. This is followed by an additional coordinate transform from cylindrical to spherical coordinates. This is needed so that the magnetic field and electron densities along each line of sight (corresponding to the spherical radial direction) can be obtained. We use the HealPIX NSIDE = 512 and the LiteBIRD frequency of 140 GHz to simulate the maps. We mask 20% of the sky map, corresponding to latitudes around the galactic plane. We use a domain size of 150 pc for the profiles [65, 66, 67, 68, 69, 70, 71]. The maps are smoothed to the beam resolution correponding to this frequency channel and are shown in Fig. 1 for $m_{a}=5\times 10^{-14}$ eV, in Fig. 2 for $m_{a}=10^{-14}$ eV, and in Fig. 3 for $m_{a}=5\times 10^{-15}$ eV. Generally, the ALP signal increases with decreasing ALP mass [43]. Also, the features of the map change with the mass of ALPs, due to changing magnetic field and electron density at the resonant locations. The high mass ALPs form nearer to the galactic center than do low mass ALPs, and that can be seen from the position of the ring-like features above and below the midplane. The north-south asymmetry is mainly a result of the magnetic field strength asymmetry, with higher signal from the southern hemisphere owing to stronger magnetic fields in the southern hemisphere [51, 52, 53, 13]. Additionally, there are local features near the ring-like features: cold spots are to the right part of the rings in the northern hemisphere, and to the left half in the southern hemisphere. This can be explained as a result of the toroidal magnetic field about the galactic center, which, after coordinate transforming to the solar system coordinates, results in a decrease in the transverse magnetic field along the line of sight.
@@ 3.2.1 Latitude-based modelling of the ALP signal at the Galactic center
Template-4 $60.5\,d\sin{\theta}+0.5\,d\exp[-(\theta-\pi/2)^{2}/2\sigma^{2}],\,\,\sigma=\pi/$ 6
@@ 3.2.2 Modeling the signal at the solar system
We form nonoverlapping slices of latitudes of small widths $\Delta\theta$ around $\theta$ in the galactic coordinate frame. The effective radial distance $r$ of the signal in the galactic center frame is calculated for a given $f$ using Eq. (3.1). Shifting to the solar system frame will locate the signal in the solar system coordinates. The net modeled signal is calculated using Eq. (3.2) and mapped to points in the solar system frame. The points that lie within the same latitude slice with mean $\theta$ in the galactic center coordinates are modeled with the same parameters $A(\theta)$ and $B(\theta)$, where $\theta$ denotes the latitude corresponding to the solar system coordinates. A template is thus obtained, which for a pixelized map may be missing some values, which can be obtained by interpolation of the signal pixels on the map. This method is followed to generate the polarized intensity template map.
@@ 4 Application of the technique using LiteBIRD instrument specification
Frequency Noise (in $\mu$K-arcmin) Beam (FWHM in arcmin)
@@
We model the simulated maps for various ALP masses ($5\times 10^{-14}$ eV, $10^{-14}$ eV and $5\times 10^{-15}$ eV) using different functions $f(d,\theta)$, with various values of the length scale $d$. For each case, we use the latitude width of $\Delta\theta=0.05$ rad in the galactic center coordinates. We vary $d$ from 12 to 20.5 kpc in steps of 0.5 kpc. The functions $f(d,\theta)$ we use are shown in Table 2. The functions are chosen such that they tend to 0 at the poles, corresponding to $\theta=0$ and $\theta=\pi$. These give us a number of template maps for each simulated map. We compare the template maps using various functions in Appendix A. Since there can be a number of resonances along a line of sight, the models capture the variation in the effect resonant location at a certain latitude $\theta$. The maximum distance of this variation is taken along the galactic plane and (at $\theta=\pi/2$) corresponds to the value $d$. Template-1 provides a slow variation of the effective resonant location with latitude away from the galactic plane, followed by a steep one, while Template-3 provides a steep variation, followed by a slow one. Template-2 and Template-4 lie between the two extremes, with Template-4 being the average of Template-1 and Template-2. We model our maps at the LiteBIRD frequency of 140 GHz and set a flux threshold of $10^{-3}\,\mu$K, which implies that we do not consider pixels below this flux threshold value in our analysis. We model the signal at different frequencies using the same spatial template, but with different spectral behavior as described in Sec. 2.
@@ 5 Conclusion
The ALP distortion signal from our galaxy due to CMB photon-ALP resonant conversion depends on the electron density and magnetic field profiles in our galaxy. The galactic halo and disk consist of ionized gases and free-electron densities. The magnetic fields in this ionized plasma are of the order of $\mu$G and determine the plasma hydrodynamics [51, 52, 53, 13]. The electron densities and magnetic field profiles in our galaxy are characterized by small-scale turbulence. This makes it difficult to make precise measurements of these profiles. The ALP signal shows non-Gaussian features [55], so it is difficult to probe it using only the power spectrum information of the CMB. In addition, simulating the ALP signal for these profiles is computationally heavy and calls for the need for faster methods to search for the ALP distortion signal.
@@ Acknowledgments
[…]

### v2

@@ 1 Introduction
Axions and axion-like particles (ALPs) are hypothetical particles predicted by the Beyond Standard Model (BSM) theories [23, 24, 25, 26, 27]. Peccei and Quinn proposed a solution to the strong-CP problem, which was identified to give rise to hypothetical particles named axions [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Axions are produced as a result of the global Peccei-Quinn U(1) symmetry breaking by the vacuum expectation value of a scalar field [32, 28, 37]. Since axions and ALPs possess mass, they are Pseudo Nambu Goldstone Bosons (PNGBs) and make one of the dark matter candidates alongside Primordial Black Holes (PBHs), MAssive Compact Halo Objects (MACHOs), Weakly Interacting Massive Particles (WIMPs), sterile neutrinos, etc., and may only weakly interact with photons [23, 24, 25, 26, 27, 41, 36, 42, 43].
@@
[…]
@@ 2.2 The spatial distribution of the electron density and magnetic field in the Milky Way
Here $\mathrm{sech}(x)$ refers to the hyperbolic secant function ($\mathrm{sech}(x)=2[e^{x}+e^{-x}]^{-1}$) which achieves the maximum value of 1 at $x=0$ and tends to 0 at asymptotic values of $x\rightarrow\pm\infty$. The $\mathrm{sech^{2}}$ form is used as it facilitates a rounder variation of the electron density at $z=0$, as compared to an exponential dependence. The $U(r)$ denotes the step function (=1=1 for $r>0$, and =0=0 for $r<0$), and the values of the parameters for the two components are shown in Table 0(b).
@@ 2.3 Simulating the ALP signal in the Milky Way
Using the formalism describe above, we can model the photon to ALPs conversion signal for different ALP masses and coupling strengths. We show a few representative simulations of the polarization maps for ALP masses $5\times 10^{-14}$ eV, $10^{-14}$ eV and $5\times 10^{-15}$ eV using the mentioned magnetic field and electron density models and coordinate transforming them to spherical coordinates, with the solar system as the origin. The observation point is the solar system, which is at the distance of about 8.5 kpc from the galactic center. To obtain the signal from different directions at this observation point, we coordinate transform to the solar system frame which is the origin. This is followed by an additional coordinate transform from cylindrical to spherical coordinates. This is needed so that the magnetic field and electron densities along each line of sight (corresponding to the spherical radial direction) can be obtained. We use the HealPIX NSIDE = 512 and the LiteBIRD frequency of 140 GHz to simulate the maps. We mask 20% of the sky map, corresponding to latitudes around the galactic plane. We use a domain size of 150 pc for the profiles [77, 78, 79, 80, 81, 82, 83]. The maps are smoothed to the beam resolution corresponding to this frequency channel and are shown in Fig. 1 for $m_{a}=5\times 10^{-14}$ eV, in Fig. 2 for $m_{a}=10^{-14}$ eV, and in Fig. 3 for $m_{a}=5\times 10^{-15}$ eV. Generally, the ALP signal increases with decreasing ALP mass [55]. Also, the features of the map change with the mass of ALPs, due to changing magnetic field and electron density at the resonant locations. The high mass ALPs form nearer to the galactic center than do low mass ALPs, and that can be seen from the position of the ring-like features above and below the midplane. The north-south asymmetry is mainly a result of the magnetic field strength asymmetry, with higher signal from the southern hemisphere owing to stronger magnetic fields in the southern hemisphere [63, 64, 65, 13]. Additionally, there are local features near the ring-like features: cold spots are to the right part of the rings in the northern hemisphere, and to the left half in the southern hemisphere. This can be explained as a result of the toroidal magnetic field about the galactic center, which, after coordinate transforming to the solar system coordinates, results in a decrease in the transverse magnetic field along the line of sight.
@@ 3.2.1 Latitude-based modelling of the ALP signal at the Galactic center
Template-4 $0.5\,d\sin{\theta}+0.5\,d\exp[-(\theta-\pi/2)^{2}/2\sigma^{2}],\,\,\sigma=\pi/6$
@@ 3.2.2 Modeling the signal at the solar system
We form non-overlapping slices of latitudes of small widths $\Delta\theta$ around $\theta$ in the galactic coordinate frame. The effective radial distance $r$ of the signal in the galactic center frame is calculated for a given $f$ using Eq. (3.1). Shifting to the solar system frame will locate the signal in the solar system coordinates. The net modeled signal is calculated using Eq. (3.2) and mapped to points in the solar system frame. The points that lie within the same latitude slice with mean $\theta$ in the galactic center coordinates are modeled with the same parameters $A(\theta)$ and $B(\theta)$, where $\theta$ denotes the latitude corresponding to the solar system coordinates. A template is thus obtained, which for a pixelized map may be missing some values, which can be obtained by interpolation of the signal pixels on the map. This method is followed to generate the polarized intensity template map.
@@ 4 Application of the technique using LiteBIRD instrument specification
[…]
@@
[…]
@@ 5 Conclusion
[…]
@@ Acknowledgments
This work is a part of the $\langle\texttt{data|theory}\rangle$ Universe-Lab, supported by the TIFR and the Department of Atomic Energy, Government of India. The authors express their gratitude to the $\langle\texttt{data|theory}\rangle$ Universe-Lab and the TIFR CCHPC facility for meeting the computational needs. Also, the following packages were used for this work: Astropy [116, 117, 118], , NumPy [119] CAMB [120], SciPy [98], SymPy [121], Matplotlib [122], emcee [112], HEALPix (Hierarchical Equal Area isoLatitude Pixelation of a sphere) [^1] Link to the HEALPix website http://healpix.sf.net[123, 124] and PySM [99].